OpenStudy (rebeccapink6398):

Convert 5.6 to a percent. A. 0.056% B. 5.6% C. 560% D. 5600% @Bella5 could u help? :D

OpenStudy (ranga):

To convert decimal to percentage, multiply by 100. To convert percentage to decimal, divide by 100.

OpenStudy (rebeccapink6398):

that didn't help! :(

OpenStudy (ranga):

To convert decimal to percentage, multiply by 100. 5.6 is a decimal. To convert it to percent, multiply by 100: 5.6 x 100 = ?

OpenStudy (rebeccapink6398):

OH! lol i got confused so... =560?

OpenStudy (ranga):

Yes, 560%

OpenStudy (rebeccapink6398):

THX! could you also help me with this>Which fraction is equivalent to 56%? thanks :D

OpenStudy (ranga):

To convert percentage to a fraction or to a decimal, divide by 100. 56% = 56 / 100. Simplify. 56/100 = 28/50 = 14/25

OpenStudy (rebeccapink6398):

THX! half of the questions was about this! lol so i got to do them! THX!!!

OpenStudy (ranga):

You are welcome.